OpenStudy (anonymous):

A cylinder and a sphere both have a radius r. What height h makes the cylinder equal in volume to the sphere? a. r b. π c. πr2 d.

hartnn (hartnn):

Volume of cylinder =\( \pi r^2 h\) Volume of sphere =\((4/3) \pi r^3\) Equate these volume equations and get h.

OpenStudy (anonymous):

d 4/3 r .........i am thinking thst it

hartnn (hartnn):

thats correct :)